OpenStudy (anonymous):

Solve for x: –4(3x – 2) = 6x + 2

OpenStudy (anonymous):

–4(3x – 2) = 6x + 2 -12x+8 = 6x +2 -12x-6x + 8 - 8 =6x - 6x + 2-8 -18x = -6 x = 3